OpenStudy (anonymous):

Write the numeral as a Roman numeral. 1654

jimthompson5910 (jim_thompson5910):

What's 1000 as a roman numeral?

OpenStudy (anonymous):

m ?

OpenStudy (anonymous):

MDCLIV

OpenStudy (anonymous):

i think.. O.O

jimthompson5910 (jim_thompson5910):

yes MathFail101 has it

OpenStudy (anonymous):

oh ya >:)

OpenStudy (anonymous):

haa thaanks guys <3

OpenStudy (anonymous):

welocome! ^_^

jimthompson5910 (jim_thompson5910):

M = 1000 D = 500 C = 100 L = 50 I = 1 V = 5 So 1654 = 1000 + 500 + 100 + 50 + 4 = MDCLIV Note: the only quirk here is that 4 = 5 - 1 = IV (ie it's like a subtraction problem instead of an addition problem)